# RealitySummary: On-Demand Mixed Reality Document Enhancement using Large Language Models

### v1

@@ 6.3. Findings
Possible Future Improvements
A desired feature among participants was the system’s ability to recognize not only text but also images. For instance, W1 faced limitations when attempting to inquire about a diagram on a page, which the system could not recognize. Similarly, they anticipated broader usage scenarios beyond text recognition, such as generating cooking recipes from fridge ingredients or providing information on real-world objects. We anticipate that enhancing the system with multi-modal LLM capabilities could address such needs. Additionally, there was interest in a “chat history” feature to recall previous interactions with the system, highlighting a need for temporal relevance in user-system interactions. Finally, participants suggested they would be more inclined to use such a system long-term if hardware constraints were overcome.
7. Limitations and Future Work
Our experiences building RealitySummary and the insights from both our formative workshop and study highlight a range of promising research directions for MR document enhancement.
Technical Limitations
We first recognize that our system exhibits numerous technical challenges that need addressing in future enhancements. Concerning AR tracking difficulties, our evaluation found that AR tracking becomes unstable with text-only documents. This instability is exacerbated in low ambient light, diminishing the tracking’s effectiveness and reliability. Such issues raise questions about the AR technology’s adaptability and dependability for daily use. Additionally, OCR bottlenecks and image tracking errors introduced complications related to processing speeds, data transmission efficiency, and reliance on the quality of images. Given the variance in document quality, the performance of OCR and image tracking can fluctuate, leading to possible inaccuracies and delays for different types of documents. Lastly, LLM generated errors show areas for improvement, particularly in situations where the text is ambiguous or highly specialized.
From Page to Documents
Currently, the system cannot display summaries, annotations, or relationships spanning multiple pages. As the system’s inputs are limited to visible pages, content not yet encountered cannot be referenced. Future research might explore a hybrid approach. For example, if a digital version of the document is available online, the system could access that content to improve summarization. These enhancements could include the types of navigation interactions (such as dynamic tables of contents or visual reading histories) highlighted by participants in our formative study. Connecting extracted text with complete digital versions of documents could enable new visual search, preview, and comparison tools that build upon the interactions seen in prior prototypes like HoloDoc (Li et al., 2019). By merging the capabilities of augmented reality with extensive digital information repositories, we believe this gap can be bridged in the future.
Beyond Text-Based Summaries
While our current work has focused on examining text-based summarizations using a large language model, our formative study (Section 3.2) also highlights the potential for visual summarization approaches, including automated mind-mapping and charting tools. Techniques like these could support more efficient search, comparison, and integration, building upon word-scale (Goffin et al., 2014) and micro visualization techniques (Parnow and Dörk, 2015) designed to reveal data in and around existing text. Integrated paragraph-level summary visualizations, such as those introduced in Caragraphs (Masson et al., 2023), are especially promising. They would be well-suited to the kinds of placements around documents that mixed reality reading tools enable.
Multimodal Inputs
The improving multimodal input capabilities of mixed reality devices and LLMs, such as visual and audio input, eye tracking, more precise hand tracking, and richer gestural interactions, create additional opportunities for dynamic reading assistance. For instance, the eye-tracking capabilities of modern MR hardware could be used to estimate gaze and incorporate that information to tailor summaries according to the viewer’s current task or area of interest. If the system also captures verbal communication in addition to written text, like in RealityTalk (Liao et al., 2022), we could use this for conversation and discussion assistance. In this scenario, summaries would reflect elements of the discussion. Combining speech and physical manipulation could allow users to dynamically request, expand, adapt, or dismiss summaries and support interactions in a wider range of contexts and environments (Oviatt, 2007). Finally, the use of multimodal LLMs to receive graphical, environmental, and gestural input have the potential to broaden these systems’ applicability.
Long-Term Evaluation
While it highlights a variety of opportunities for augmented reading tools, our in-the-wild study is still a short-term evaluation. Long-term studies, in which participants interact with these systems for weeks or even months will almost certainly reveal deeper insights and challenges. Long-term studies are challenging given the current limitations of mixed reality hardware like the Hololens 2 and Vision Pro, which are not comfortable enough for everyday use. We anticipate that the continuous improvement of the headset hardware will make long-term studies feasible in the future.
@@ 8. Conclusion
RealitySummary is the first example of an on-demand MR reading assistant that uses OCR and an LLM to generate reading aids for any physical or digital document. We hope our study opens up new opportunities for on-demand and natural language-powered reading assistance and inspires the HCI community to continue exploring such systems to realize the full potential of intelligent reading tools.

### v2

@@ 5.4. Insights
Trust of AI-Generated Content
Many participants approached the system’s responses with caution, often verifying the answers with the reading material. They highlighted the necessity of consulting additional sources for in-depth or technical readings when depending on the system. There was a general agreement among participants about their reluctance to fully trust an always-on reading assistant in their daily activities. Conversely, we noted some intriguing discussions regarding how participants perceive our system, either as a mere tool or more like a personal companion. While most participants engaged with the system as a conventional chatbot, some regarded it more as a personal companion than a mere tool. Notably, W3 and W5 described the system as a friend for discussing and inquiring about their readings. The system’s ability to provide always-on, immediate responses facilitates this personification.
5.5. Limitations
Overall, the system design of the second prototype enhanced its applicability and stability for real-world use. Compared to the first prototype on the Hololens 2, the Vision Pro’s superior rendering quality and broader field of view provided a better user experience. Additionally, the simple system design greatly improved usability despite the absence of document tracking. While the second prototype showed great potential, we also identified several feature-level limitations, along with inherent study-related constraints.
Lack of History and Recall Support
While the minimalist design effectively reduces visual clutter and conserves screen space, it also limits the system’s overall utility. Although users appreciated the clean and simple interface, many expressed frustration with the lack of options to save or revisit previously generated responses. Several participants suggested that a history feature or the ability to store responses for future reference would substantially enhance the user experience.
Challenges with Screen Viewing for Extended Use
Many participants raised concerns about the difficulty of clearly viewing phone and computer screens through the Vision Pro without dramatically enlarging the text. While this worked for short-term tasks like quickly reviewing an assignment, participants found it frustrating for longer viewing periods, particularly when using laptops. Since much of their work is done on computer screens, this limitation poses a challenge for assessing the system’s effectiveness in supporting realistic, work-related tasks.
Need for a Longer-Term Study
Our in-the-wild study was inherently limited by its short-term, public use — with each session lasting only one hour per participant. As a result, participants did not have the opportunity to fully engage with the system in their personal workflows. For instance, due to privacy concerns related to screen recording, participants rarely used our system for more personalized content, such as emails, messages, or calendars. Furthermore, since participants were encouraged to explore various scenarios, they tended to experiment broadly rather than focusing on any specific task in depth. While the in-the-wild study provided valuable insights into real-world use cases, it only captured short-term interactions in public spaces. There is still much to learn about the sustained, regular use of reading assistants in daily life, particularly in how users interact with information over time, revisit previous queries, or support ongoing tasks.
6. Study 3 RS-Diary: Diary Study with the Third Prototype
To explore the long-term use cases, we developed a third iteration, incorporating feedback from the in-the-wild study. This version introduced a multiple responses feature and added the ability to view or mirror the content on an external device, addressing the readability issues for the long-term use of the Vision Pro.
Figure 9. Snapshots captured from the Vision Pro during our diary study.
6.1. Implementation
The final version of RealitySummary builds on the earlier Vision Pro implementation with two key improvements: 1) Users can now add multiple draggable cards on-demand by either making a new query or clicking an “add” button. User can spatially anchor these cards anywhere in their immediate environment and can also re-select cards to update them with new prompts or text from the scene. These cards are stored as JSON objects in a real-time database. 2) The system allows users to mirror their device screen within the Vision Pro for high-resolution viewing. This functionality — implemented through a WebRTC video call or via Mac Virtual Display on the Vision Pro — lets users view their external devices at full resolution in the MR environment and also allows RealitySummary to more accurately capture and process on-screen content.
6.2. Method
Five authors (A1-A5) used the system over the course of 3 weeks, integrating it into their day-to-day activities and accumulating 30 total hours of use. Due to physical fatigue and battery limitations of the Vision Pro, each session lasted less than 2 hours. The authors independently documented their use cases, insights, frustrations, thoughts, and reflections while using RealitySummary, which we later thematically coded. During the study, each participating author recorded their observations via a phone microphone, following a think-aloud process. Additionally, the first author conducted follow-up interviews to gain deeper insights into their experiences.
6.3. Results
Throughout the diary study, participants used the system in various settings, including labs, kitchens, outdoor areas, and offices, though the majority of usage took place at desks in work environments. They engaged with a wide range of reading materials, such as academic papers, news articles, code, emails, calendars, Google Scholar, classroom assignments, lecture slides, textbooks, and YouTube videos. Participants quickly adapted the system to fit a variety of personal, academic, and technical tasks. For example, A4 used the assistant not only to summarize email content but also to gather higher-level insights, such as identifying how many emails were sent by a particular sender. A2 utilized the system to support programming assignments, where they used the system to break down complex coding problems into manageable steps and provided detailed explanations of specific code snippets. A2 also leveraged the system for learning aids, such as generating to-do lists and quiz questions based on problem descriptions. The ability to spatially organize tasks and concepts allowed him to effectively manage his workload, maintain focus, and dynamically track his progress. Additionally, in his language learning efforts, A2 used the assistant to generate contextual vocabulary lists from Japanese reading materials, which helped him focus on the most relevant and frequently used words. A5 employed it to extract a step-by-step sequence of keyboard commands from an article on rebooting a MacBook. Overall, the diary study highlights the system’s flexibility in supporting diverse tasks, from email management and coding to technical troubleshooting, language learning, and academic research.
6.4. Insights
From these experiences, participants shared various insights, frustrations, and potential future improvements for the current prototype. Our diary study also provided insight into the use of an always-on reading assistant in more realistic work-oriented settings.
Quick and Simple Summary Responses Meet Many Immediate Needs
Participants valued the system’s ability to deliver quick, concise responses, which were often sufficient for immediate clarifications. A2, for instance, appreciated how the system provided brief answers when they needed a rapid comparison of philosophical concepts, like distinguishing between act and rule utilitarianism. In many cases, just a high-level summary captured through manual input was enough to satisfy their needs. A4 found this method to be a useful shortcut, as pressing a button was faster and less disruptive than speaking and often provided exactly the information they were seeking. These simple, proactive summaries were particularly beneficial for tasks that required quick comprehension and minimal interruption.
Minimal Context Switching Enhances Focus and Flow
Participants appreciated the integration of LLM into the MR interfaces, in part, because it helped them minimize context switching, allowing them to maintain greater focus. The system’s automatic text extraction feature enabled participants to stay in a flow state. For example, A2 noted that the interface allowed them to concentrate on their work without the mental disruption of switching between tasks or manually inputting context, even when using a computer screen. This seamless workflow kept participants engaged, enhancing productivity and reducing cognitive load. A1 highlighted how the system functioned like a “teacher” by offering real-time assistance without requiring them to step away from their tasks. Participants also remarked that on-demand support boosted their efficiency during complex tasks like coding or writing, where focus is critical.
Using Generated Cards as Peripheral References
Participants highlighted the advantages of peripheral output modalities. For instance, A2 used the system as a reading aid while learning Japanese. Instead of frequently opening a dictionary, which disrupted the reading flow, they generated vocabulary flashcards around the document. This provided the needed language support without interrupting the reading and learning process. A4 also noted that they would often cluster and group these generated cards thematically to the sides of the display for quick reference and then delete them once they were no longer needed. Similarly, A1 used the system to “pin” generated code versions and key concepts around their workspace, allowing for easy access without breaking focus and enabling them to quickly revisit important points as needed.
Spatial Organization Improves Recall and Multitasking
A key finding from the study was how participants leveraged the system’s spatial organization features, similar to arranging post-it notes on a board. Both A1 and A2 utilized this capability to group related concepts together. For example, A2 clustered cards on CPU architecture, placing related topics like buses and registers nearby, much like organizing post-its by theme. This spatial arrangement enabled quick access and improved recall, as users could visually track where specific information was placed. Similarly, A1 pinned cards documenting different problem-solving methods across their workspace, ensuring easy access when needed. This post-it-like organization supported memory retention and task management by allowing users to arrange information in a way that aligned with their thought processes.
6.5. Limitations
Challenges Controlling What to Capture
Currently, the system captures the user’s entire field of view, but participants expressed a desire for more control over what is captured. For example, A1 struggled to get the system to ignore personal emails displayed next to the article he was reading, leading to summaries that inadvertently included private information. Similarly, A2 wanted to focus on a specific code snippet to obtain more accurate results while debugging. Additionally, the lack of visual feedback left users uncertain about what OCR content or context was being processed in each request. To address these issues, the system could offer enhanced visual feedback and allow users to control specific areas of the screen, such as through gaze interaction, which would help mitigate privacy concerns and prevent the inclusion of irrelevant or personal content.
Cumbersome Re-Prompting Process for Common and Repeated Tasks
Participants frequently found themselves re-prompting similar queries for routine tasks, such as requesting bullet point summaries or filtering by specific criteria. A potential improvement could be the addition of dynamic buttons that save commonly used prompt types, allowing users to easily reuse them later. Additionally, participants expressed frustration with testing out lots of different prompts to generate a complex output, such as aggregating the Google Scholar page into a timeline. Saving or sharing the same specific instructions for repeated tasks might mitigate these problems.
Managing Growing Numbers of Summaries
In the first prototype, participants found it distracting to have multiple summary cards anchored to the image in their workspace. However, they did not find this frustrating when they had control over the placement and could easily delete the cards. Additionally, participants noted that the system could sometimes generate an overwhelming number of cards. For instance, A2 mentioned that while analyzing a poem, the system created a card for each sentence and its meaning, making it difficult to manage information overload. This feedback suggests that there is potential value in exploring new interactive and semi-automated methods for efficiently sorting, snapping, organizing, filtering, and removing these cards. Such improvements could enhance the usability of the system and help users better manage the content generated.
7. Lessons Learned from Three Studies
Drawing on results from across our prototyping and evaluations, we close by showcasing the unique benefits of combining MR and AI tools, enumerating potential use cases, and highlighting opportunities for future research.
7.1. Unique Benefits of Combining MR and AI
We identified two unique benefits of combining MR with AI: 1) implicit input and 2) peripheral output.
Always-On Implicit Input
The always-on camera in MR enables implicit input, distinguishing it from the explicit input required in desktop interfaces, such as opening an app, typing queries, or copy-pasting contextual text. Participants in in-the-wild and diary studies appreciated the system’s ability to capture information without requiring explicit effort. This always-on input allows for hands-free exploration in both spatial and temporal contexts, such as scanning entire bookshelves as a query or utilizing the history of flipped pages. Implicit input reduces cognitive load and minimizes context switching, as users are not required to actively specify what information should be used for the LLM’s context.
Peripheral and Background Output
We found that MR’s spatial output allows information to be presented in a peripheral, background manner. In contrast, other output modalities, such as audio or screen-based displays, often require the user’s foreground attention to consume generated content, disrupting focus and interrupting workflow. With MR and peripheral outputs, however, AI-generated content becomes less intrusive, as it can be easily ignored if not needed, simply by choosing not to look at it. For example, participants in the diary study often appreciated this feature for learning aids, such as understanding textbooks or accessing always-on vocabulary references when learning a foreign language. This peripheral output is a unique aspect of the integration of MR and AI. We believe this capability would promote more calm human-AI interactions in the future, reducing disruptions to focused attention by combining implicit input with unobtrusive background support.
7.2. Emergent Use Cases for Mixed Reality Text Augmentation
Both the in-the-wild and diary studies revealed a variety of interesting emergent use cases. To provide a more comprehensive view of how users engage with these systems, we have categorized these into seven groups:
Summarizing Content
The primary use case was to summarize content. This was the most commonly observed behavior across various contexts. Participants summarized a wide range of texts, including research papers, blog posts, news articles, books, and instructional guides.
Extracting Information
Participants also utilized the system to extract key information, which can be translated into a more consumable format. This aligns with design elicitation results, which suggest tables of contents, keyword lists, timelines, and comparison tables. For example, one participant extracted vocabulary flashcards to assist with foreign language reading, while another participant created a table of keyboard shortcut commands from a troubleshooting guide to use as a reference. In these examples, the consumable format, such as tables or vocabulary lists, serves as helpful references for their tasks.
Querying External Sources
Participants frequently used extracted text as a query to retrieve additional information from external sources. Echoing the information types identified in our elicitation study, participants queried specific names or keywords from their reading materials. For instance, one participant retrieved details about a player, team, and standings from a sports-related article, while another fact-checked claims on-demand when reading news articles.
Asking for Suggestions
Beyond the above use cases identified in the elicitation study, participants frequently asked for recommendations or suggestions based on the current context. For instance, users asked for book recommendations when looking at a bookshelf. Beyond the reading context, one participant asked for advice on which trash bin can be used based on disposal instructions, while others inquired about healthier menu options or products. They also asked for recommendations from external sources. For instance, participants asked for suggestions for relevant research papers when reading academic articles.
Aggregating Information
An interesting use case involved using the system to aggregate high-level information. Unlike simple information extraction, aggregation often involves operations like filtering and counting. For example, some participants used the system to count emails based on specific filters, while others filtered and counted first-author papers from a researcher’s Google Scholar page. This allowed users to perform quick, approximate queries that would normally require more complex searches.
Transforming Content
Participants also used the system to transform their reading content into interactive, actionable outputs. For instance, some participants applied our system to generate quizzes from class notes or textbooks, while others decomposed high-level assignments into step-by-step to-do lists. One participant asked the system to create reminder cards that remained spatially persistent. Participants also use our system to translate from one language to another. Unlike information extraction, these tasks involved converting content into actionable, interactive outputs.
Generating New Content
Finally, participants generated new content based on existing text. In one example, a participant requested multiple variations of a written paragraph, which they then organized spatially. Others used the system to create code based on instructions or to debug existing code snippets.
These emergent use cases highlight novel ways of interacting with AI-generated information, allowing users to engage with individual snippets and convert them into dynamic outputs. They also illustrate how participants adapted the assistant to handle different types of content, which suggests a broad space of possible designs for such systems.
7.3. Future Opportunities and Design Implications
As both MR and LLM technologies continue to evolve rapidly, considerable opportunities exist for tools that more richly integrate them. Based on our three studies, we identify several key opportunities for future research and design.
Enabling Greater Control Over Capture
While the system’s implicit input offers significant benefits, it also raises privacy concerns due to the automatic capturing without the user’s constant awareness. To address this, future designs should give users control over what data is captured. As noted in the diary study, participants expressed a desire for more transparency and control over the captured content. One approach could involve providing visual feedback, allowing users to see and select the specific areas or information being captured. Future systems should balance the benefits of implicit input with privacy considerations by offering user control and visual feedback.
Improving Organization and Filtering of Interactive Cards
Participants expressed the need for more efficient ways to manage the growing number of spatially organized content cards. While some diary study participants manually arranged these cards around documents, others suggested more advanced organizational tools. Features like sorting, snapping, filtering, or automatically organizing cards could simplify workflows and reduce clutter. Such capabilities would be particularly beneficial for long-term tasks, where the accumulation of content can become overwhelming. Integrating these organizational tools would enhance usability, supporting users in managing dynamic and task-specific information over time, which enables better workspace management for daily tasks.
Balancing Between Reusable Prompts and On-Demand Custom Queries
Our studies highlighted the importance of accommodating both reusable prompts and custom queries. Each has its advantages: prepared prompts can generate complex outputs such as timelines or comparison tables, while custom queries offer users flexibility in generating content. The diary study participants expressed a desire for dynamic interfaces that could quickly reuse and adapt prompts to different contexts. For example, some suggested having customizable buttons that automatically generate summaries or lists in a consistent style. Others mentioned the need for a more efficient way to reuse queries across tasks, such as repeatedly generating lists or comparisons for class topics. This feedback highlights the importance of designing adaptive interfaces that allow users to seamlessly reuse and modify their prompts, improving efficiency and user satisfaction.
Facilitating Seamless Transitions Between Screen and Space
Participants also valued the ability to drag content cards from the system into their computer, such as pasting generated answers directly into coding tasks. Some participants also envisioned this feature being useful for writing tasks, allowing for a smooth transfer of generated content. Future systems should explore leveraging cross-device interactions between MR and other devices to facilitate seamless transitions, enhancing the workflow between screen-based and spatial interactions.
Expanding Beyond Text Input to Capture Images and Everyday Objects
Participants expressed a strong desire for the system to recognize not only text but also images. For example, the in-the-wild study revealed the limitations of inquiring about a diagram that the system could not recognize. Users also anticipated broader applications beyond text recognition, such as generating recipes based on ingredients in the fridge or providing information on real-world objects. Enhancing the system with multi-modal LLM capabilities could address these needs, expanding its utility beyond text-based interactions.
Personalization Through Always-On History
Lastly, participants showed interest in an “always-on” history feature that would allow them to recall previous interactions with the system, creating a more personalized experience. They speculated that such a system could eventually serve as a personalized assistant, continuously capturing and understanding context through an always-on camera. Although current hardware limitations, such as battery life and physical fatigue, prevent continuous long-term use, we believe future MR interfaces will seamlessly integrate with AI to provide both context-aware and personalized assistance based on users’ daily activity histories.
@@ 8. Conclusion
RealitySummary is the first on-demand MR reading assistant that combines OCR and LLM technologies to generate spatially-situated on-demand reading aids for both physical and digital documents. Through our empirical studies, we aimed to highlight new possibilities for on-demand, AI-powered reading support. This paper presented an iterative development of the system, along with key findings and lessons learned from three studies. We hope our work will inspire the HCI community to further explore the potential of integrating MR and AI in future applications.